### 07 Feb 2024 v1

# Biological computation through recurrence

Abstract
One of the defining features of living systems is their adaptability to changing environmental conditions. This requires organisms to extract temporal and spatial features of their environment, and use that information to compute the appropriate response. In the last two decades, a growing body or work, mainly coming from the machine learning and computational neuroscience fields, has shown that such complex information processing can be performed by recurrent networks. In those networks, temporal computations emerge from the interaction between incoming stimuli and the internal dynamic state of the network. In this article we review our current understanding of how recurrent networks can be used by biological systems, from cells to brains, for complex information processing. Rather than focusing on sophisticated, artificial recurrent architectures such as long short-term memory (LSTM) networks, here we concentrate on simpler network structures and learning algorithms that can be expected to have been found by evolution. We also review studies showing evidence of naturally occurring recurrent networks in living organisms. Lastly, we discuss some relevant evolutionary aspects concerning the emergence of this natural computation paradigm.
I Introduction
The survival of any living system depends on its capacity to sense multiple signals in its surroundings, integrate those external signals, and activate an adequate response. Organisms need to derive information from both the spatial and temporal structure of these signals [1]. Temporal information, in particular, is crucial to anticipate changes in the environment. In these cases, having some memory mechanism is clearly beneficial. In the mammalian brain, for instance, processing the temporal structure of visual stimuli gives information on the direction and speed of objects, as well as on the duration of and time between events [2]. In the somatosensory system, the contribution of temporal encoding is not limited to dynamic aspects of the stimulus, such as motion detection, but is also involved in stationary aspects such as object and texture discrimination [3]. It has been postulated that recurrence due to extensive feedback can underlie the temporal information-processing abilities of cortical neural networks [4, 2].
The importance of processing temporal information is not limited to neural systems. Individual, non-neural cells such as bacteria are able to relate groups of events that tend to occur together or in a specific order, as a form of associative memory [5]. For example, Escherichia coli prepares itself for low oxygen conditions when it detects a rise in temperature, an indication that it has been ingested by a mammal [6]. Similarly, Candida albicans cells, a fungal pathogen, upregulates oxidative stress resistance genes upon detection of an increase in glucose levels (which indicates they have entered a host), preemptively reacting to its immune system [7].
Apart from this associative memory capacity, recent studies indicate that microorganisms display both short- and long-term memory. For instance, the stress response of Bacillus subtilis relies not solely on its present growth conditions but also on its past [8]. The way in which this record of previous history —i.e. memory— is integrated and stored in bacteria cells is the subject of much interest. It has been postulated that this adaptability relies in general on a complex network of interactions between genes and proteins, frequently at the level of transcriptional regulation [9, 10]. These types of networks are called gene regulatory networks.
Gene regulatory networks from multiple organisms have been found to possess a structured recurrent architecture reminiscent of the one underlying a computational paradigm known as reservoir computing (RC) [11]. In this scheme, temporal computations emerge from the inherent structure and dynamics of the network a very efficient learning process [12, 13]. This paradigm was also proposed to underlie brain computations under the name of liquid-state machines (LSMs) [14].
The key characteristic of RC is that it separates memory encoding and prediction in different network substructures, as depicted in Fig. 1.
Figure 1: Structural and functional organization of the gene regulatory network of Escherichia coli. The genes represented in green form a recurrent subgraph with cyclic paths (the reservoir) that can maintain a record of the recent history in its dynamics. The genes in red constitute a directed acyclic subgraph (the readout) that interprets the information encoded in the reservoir state. The network architecture shown here was obtained from the Ecocyc database [15]. The representation is adapted from Gabalda-Sagarra, Carey, and Garcia-Ojalvo [11].
First, a substructure with recurrent connections is needed to encode history: the reservoir. The reservoir serves two functions simultaneously:
  1. 1.
    It acts as a nonlinear high-dimensional expansion of the input signal(s).
  2. 2.
    It allows the network to retain information for a certain time, giving memory to the system.
The second substructure, the readout, is a feedforward architecture located downstream of the reservoir (see red nodes in Fig. 1). The readout layer uses the history record encoded in the state of the reservoir, combined with the current input, to make a prediction or classification. This separation of roles allows to focus the training process (i.e. the modification of the network’s synaptic strengths so that the desired output is produced) solely on the readout [16].
In the classical RC approach, also known as an echo-state network (ESN), the reservoir is generated randomly, and only the readout is trained. Since feedforward architectures are much easier to train than recurrent ones, the RC framework has the computational power of a recurrent network and the ease of training of a feedforward one [2]. RC has demonstrated outstanding performance across various benchmark tasks [17, 18, 19].
The simplicity of the network architecture underlying RC, and the evidence that natural gene regulatory networks exhibit this structure, support the expectation that recurrent biological networks might underlie the notable information-processing capabilities of living systems. In the following section we introduce a basic mathematical description of how recurrent networks can compute. Next we discuss the potential application of this paradigm in biological systems. Finally we look into the evolutionary aspects of this natural computation framework.
II How recurrent biological networks compute
II.1 Basic description
In what follows, we focus on gene regulatory networks as information-processing substrates, although the same mathematical description can be applied to other biological networks (such as neuronal networks, which in fact are the ones most commonly considered in the literature within this context). The dynamics of a gene regulatory network can be represented mathematically using models similar to those typically employed to describe artificial neural networks (ANNs). Specifically, the expression level of a set of $N$ genes at a given time $t$ can be denoted by a vector $\mathbf{x}_{t}\in\mathbb{R}^{N}$, where $t=1\ldots T$ are a set of discrete time points at which the network activity is being monitored. The expression of a gene depends on the levels of the genes that regulate it, and potentially on other inputs $\mathbf{u}_{t}\in\mathbb{R}^{N_{\mathrm{in}}}$ coming from outside the network. This can be represented by the following update rule:
$\mathbf{x}_{t}=\mathbf{f}(W\mathbf{x}_{t-1}+W^{in}\mathbf{u}_{t}),$ (1)
where $\mathbf{f}$ is an element-wise activation function whose argument is a weigthed sum of the expression level of the regulating genes and the external inputs. The effect of those regulating elements is represented by the weight matrices $W\in\mathbb{R}^{N\times N}$ and $W^{in}\in\mathbb{R}^{N\times N_{in}}$, respectively. In the ANN literature, the activation function $\mathbf{f}$ typically takes for the form of a sigmoid (e.g. a hyperbolic tangent), which in fact is usually an excellent representation of the activation of a gene by its inputs [20, 21].
There is no constraint in the architecture of the network represented by the connectivity matrix $W$. In particular, there is no need for the network to be feedforward, as is commonly the case in basic ANN (including deep-network) architectures. In fact we consider here that the network underlying $W$ is recurrent, as in the ESN architectures used in RC [16]. The input signals $\mathbf{u}_{t}$ correspond to genes (or other biochemical elements) that are fully upstream of our network (orange nodes in Fig. 2).
Figure 2: An echo state network. The upstream genes (orange) send the external inputs ($\mathbf{u}_{t}$) to genes of the reservoir (green) with strengths given by the input weight matrix $W^{in}$. The reservoir genes are coupled to each other by recurrent connections defined by the matrix $W$. Finally, the readout genes (blue) read the information encoded in the reservoir to make decisions. Only the readout weights $W^{out}$ (red arrows) are trained, and this in done in such a way that the output $Y$ approximates the target output signal(s).
In turn, the network output is given by a weighted sum of the states of the reservoir genes:
$\mathbf{y}_{t}=W^{out}\mathbf{x}_{t}$ (2)
where $\mathbf{y}_{t}\in\mathbb{R}^{N_{out}}$ is a vector of all predicted outputs, and $W^{out}\in\mathbb{R}^{N_{out}\times N}$ is the weight matrix for the reservoir-readout connections, as schematized in Fig. 2.
The values of the readout weights $W^{out}$ determine the computational capabilities of the network. In the artificial neural-network context those weights are trained in a supervised manner, to ensure that the network output $Y$ reproduces a given target output in the presence of the corresponding input. This can be done very efficiently using a matrix operation such as ridge regression:
$W^{out}=Y^{\mathrm{target}}X^{T}(XX^{T}+\gamma^{2}I)^{-1},$ (3)
where $Y^{\mathrm{target}}\in\mathbb{R}^{N_{out}\times T}$ is a matrix with all expected/target outputs over time, $X\in\mathbb{R}^{N\times T}$ is a matrix with the state of the reservoir nodes over time, $X^{T}$ is the transpose of $X$, $I$ is the identity matrix, and $\gamma$ is a regularization coefficient, which helps avoid overfitting (see Sec. 8 of Lukoševičius and Jaeger [22] for details).
It is worth noting that for the ESN approach to work, the network needs to be appropriately tuned. Specifically, the recurrent reservoir should satisfy the so-called echo state property: its state $\mathbf{x}_{t}$ should be uniquely defined by the fading memory of the input $\mathbf{u}_{t}$ [17]. In other words, for a long enough input $\mathbf{u}_{t}$, the reservoir state $\mathbf{x}_{t}$ should not depend on the initial conditions that were present before the input. It is empirically observed that the echo state property is obtained for any input if the spectral radius (i.e. the maximum absolute eigenvalue of the reservoir connectivity matrix $W$) is smaller than 1. Large spectral radius values can lead to reservoirs exhibiting multiple fixed points or periodic, even chaotic (when sufficient nonlinearity in the reservoir is reached) attractors, which violate the echo state property and prevent efficient training.
II.2 Alternative training algorithms
The echo-state property described above implies that the reservoir is quiescent in the absence of external inputs. However, biological networks usually display spontaneous activity characterized by intricate and irregular patterns. In fact, recurrent networks with chaotic self-sustained dynamics have been found to be highly effective for computational purposes [23]. In particular, chaotic networks exhibit faster training times and generate outputs that are more accurate and robust compared to non-chaotic networks. Given the inability of the RC paradigm to operate in these conditions, a new training algorithm is needed. For that purpose Sussillo and Abbott [23] introduced a learning algorithm, called FORCE (First-Order Reduced and Controlled Error), that enables training the network even when it operates in a chaotic regime. Interestingly, this algorithm is more biologically realistic than the one used in ESNs [16]. Given the occurrence of self-sustained dynamics in natural networks, and the relative realism of the FORCE algorithm for potentially training those networks, we discuss it in some detail in what follows.
To be operational, the learning algorithm must induce in the network a transition from chaotic to non-chaotic dynamics during training. This is precisely what the FORCE procedure achieves. Specifically, Sussillo and Abbott [23] show that the chaotic dynamics during training can be quenched by using either external feedback loops (Fig. 3a) or internal synaptic modifications (Figs. 3b and 3c).
Figure 3: Network architectures for the FORCE algorithm [23]. In all three cases, a reservoir (large circle with green nodes) drives a linear readout unit through weights that are modified during training (red arrows coming into the red node). Only connections shown in red are subject to modification. (a) Feedback to the reservoir is provided by the readout unit. (b) Feedback to the reservoir is provided by a separate feedback network (smaller circled network). The neurons of the feedback network are recurrently connected and receive input from the reservoir through synapses that are modified during training (in red). (c) Network with no external feedback. Instead, feedback is generated within the reservoir and modified by applying FORCE learning to the synapses internal to the network (red).
All three schemes lead to the generation of intricate yet controlled outputs. Driven by any of these two strategies, the key for suppressing chaos is making strong and rapid synaptic modifications. In this regard, FORCE learning works in a manner distinct from traditional neural network training. Typically, conventional training involves a series of adjustments aimed at gradually diminishing initially large errors in the network output. In contrast, the FORCE algorithm attains a large reduction in the error’s size at the time of the first weight update, and then maintains small errors throughout the training process. In order to achieve this, the method makes use of the recursive least-squares (RLS) algorithm [24] as the update rule for the trained weights (see Sussillo and Abbott [23] for details). Finally, by the end of the training period, modification is no longer needed, and the network can generate the desired output autonomously.
Lastly, it is worth emphasizing that the FORCE algorithm allows us to train not only the readout weights of the network, such as in the original RC framework (Fig. 3a), but also RNNs with the architectures shown in Figs. 3b and 3c, in which modifications are not restricted to the readout weights, but instead include internal synaptic modifications. These architectures match biological circuits better than that of Fig. 3a, where feedback comes directly (and exclusively) from the readout unit. By enabling training of different architectures, the FORCE algorithm extends the original RC framework, and gives more flexibility and biological feasibility to the concept of computation through recurrence.
II.3 Using dynamical systems theory to understand recurrence-based computations
Since supervised training algorithms specify the computation to perform without constraining the network on how to implement this computation, the precise manner in which trained RNNs realize their target functions remains an open question. In essence, the resulting networks have predominantly been regarded as black boxes. This differs from network models deliberately designed to implement specific established mechanisms, such as those proposed, for instance, by Hopfield [25] and Wang [26]. One way of “opening the black box”, proposed by Sussillo and Barak [27], is to treat the network as a nonlinear dynamical system.
Sussillo and Barak [27] studied not only the existence of fixed points (both stable and unstable) and the linearized dynamics around them, but also the linearization in areas of phase space that are not true fixed points, but merely points in phase space through the system flows very slowly (slow points). They also present a simple optimization technique for locating such regions. Linearization around each one of the fixed and slow points provides a mechanistic description of how high-dimensional RNNs implement their computations [27]. This last point is key in order to suggest hypotheses for how the task might be solved by the biological network. Therein lies the importance of this dynamical-system based approach.
Identifying the fixed and slow points of an RNN requires optimization tools that rely on hard-coded analytic derivatives, e.g. of the hyperbolic tangent function (see e.g. Section 3.2 of Sussillo and Barak [27] for details). Algorithms to find fixed points have been developed for simple (Elman) RNNs [28], and Hopfield-like RNN architectures [27, 29], but can be inconvenient for complicated models, or for studies that involve a combination of different RNN models. In this sense, Golub and Sussillo [30] introduced an open-source Tensorflow toolbox for finding fixed points and linearized dynamics in arbitrary RNN architectures. This software profits from TensorFlow’s highly efficient automatic differentiation back-end to execute the necessary optimizations and compute Jacobians, essential for the identification and characterization of fixed points.
III Recurrent networks in biology
III.1 The mammalian cerebral cortex
Cortical circuits were inspirational to Maass, Natschläger, and Markram [14] when conceiving LSMs. It has been proposed that cortical networks are likely to operate following an RC paradigm, or at least that they incorporate most RC principles. One indication of this is the fact that neurons vary morphologically and functionally throughout the cortex. Heterogeneity in neural characteristics has been observed in cortical neurons, the retina, and the primary visual cortex, as well as in networks cultured in vitro from dissociated neurons [31, 32, 33, 34, 35, 36]. This diversity of neural components results in very heterogeneous dynamics, which can be exploited by RC-like computation. In addition, cortical networks present a highly recursive nature [37].
Nonetheless, discerning the computational basis of cortical circuits is still today an open question in neuroscience, and efforts to clarify to what end recurrence is exploited in cortical circuits are under way [31, 32]. In what follows, we review related studies on cortical regions, including prefrontal and visual areas.
Cortico-striatal system
An early RC-type model of the cortico-striatal system was proposed by Dominey, Arbib, and Joseph [38] to understand the mechanism of context-dependent oculomotor (eye movement) behavior [39]. The model consists of two sub-networks, based on the known cortical/subcortical anatomy of the primate oculomotor system. The first sub-network corresponds to the prefrontal cortex, which is associated with the planning of cognitive behavior, personality expression and decision making, besides moderating social behavior [40]. The second sub-network corresponds to the striatum, which is a subcortical area.
In the model of Dominey, Arbib, and Joseph [38], the prefrontal sub-network has fixed recurrent connections between its neurons (reservoir); while connections from the prefrontal cortex to the striatum are modifiable (readout) [41]. The neural activity in the prefrontal cortex network generates unique patterns of activity in response to the sequential visual inputs. The striatum associates these state-encoding patterns with the correct outputs, which represent the oculomotor movements.
Later, Dominey and Ramus [42] extended this model to the context of language processing in infants. Specifically, the authors showed that the previously mentioned cortico-striatal model (called a temporal recurrent network) simulates the capabilities of babies to learn the sequential and rhythmic structure of language. Indeed, the recurrent neural network is capable of representing serial and temporal structure with no pre-exposure, and through exposure these internal representations can become bound to behavioural responses. A modified version of the model, with an additional recurrent network with short-term memory, was however needed to reproduce the performance of babies in extracting abstract structure (or a hidden rule) from sequential inputs. These works illustrate how inherent representational abilities of the serial and temporal structure of language might emerge from a shared neural architecture, separate from what is needed for representing abstract structures.
The cortico-striatal model was further combined with neurophysiological models of language processing to learn grammatical constructions in sentence processing [43, 44] and with an RC framework to improve the performance of the learning algorithm [45]. The detailed history and the developments of the cortico-striatal model was reviewed by Dominey [46].
Visual cortex
The visual cortex is the part of the cerebral cortex that processes visual information. It is divided into the primary visual cortex (early visual areas) and visual association cortex (higher-order visual areas) [47]. Nikolić et al. [33] obtained experimental evidence showing that early visual areas in cats have fading memory properties, as demanded by LSMs. Specifically, in vivo spiking activities of 60–100 neurons were obtained for a sequence of different visual stimuli (letters). The authors showed that the neuronal responses contained information not only about the letter being presented but also about the preceding letters. In other words, the population activity encodes both the current and previous stimuli. In fact, the identity of the different visual stimuli (in this case letters) could be recovered with high reliability from the simultaneously recorded spike trains of neurons (corresponding to reservoir states) by a simple linear classifier emulating linear integrate-and-fire readout neurons [33].
The new perspective that the primary visual cortex can perform time-dependent computation for sequential inputs based on memory mechanisms contrasts with the conventional one that the brain executes frame-by-frame computation based on memory-less hierarchically organized feedforward architectures [48]. The limitation of the conventional viewpoint is that frame-by-frame computation cannot process dynamical visual scenes.
III.2 Gene regulatory networks as reservoirs
Biological computation through recurrence does not necessarily need to arise in neural substrates, but can be supported by other biological networks. In this section, we review two example studies involving gene regulatory networks (GRNs). The first study looks into the ability of Escherichia coli to discriminate between complex temporal patterns of chemical inputs. The second one studies the memory capacities of the GRNs of five evolutionary distant organisms.
Classification of chemical inputs through E. coli’s GRN
Jones et al. [49] proposed that the gene regulatory network (GRN) of the bacterium Escherichia coli can act as a computational reservoir, receiving input signals from a sequence of chemical stimuli and acting upon a set of target proteins as an output. Specifically, the authors exposed a liquid culture of E. coli cells to sequences of inputs that represent different combinations of temperature and chemical conditions of the environment. Subsequently they studied the transcriptional state of the GRN using microarrays, to test whether the network was capable of making complex perceptual classification of the different inputs into desired classes. Numerical simulations corresponding to the above experimental design confirmed that the proposed RC framework was partially successful in an XOR problem, and showed good performance in a separation problem.
This work motivated an experimental study in which Didovyk et al. [50] genetically engineered bacterial cell cultures with the goal of realizing a variety of simple classifiers separating different chemical inputs. The outputs were then aggregated to form a more complex classifier.
Memory encoding through recurrence in GRNs
In the previous example, the GRN used as putative reservoir does not include any recurrences other than self-regulations of some sparsely interconnected nodes, and the system was tested with a task that does not require memory (classification task). Fully recurrent topologies are necessary, however, for processing time-dependent information. In this sense, Gabalda-Sagarra, Carey, and Garcia-Ojalvo [11] studied the computational capabilities of the GRNs of five evolutionary distant organisms (from bacteria to humans) in standard memory-demanding tasks. In all cases, they recognized a cyclic recurrent structure, constituted by a small fraction of the overall number of genes, which is crucial for dynamic encoding and information integration (reservoir). Meanwhile, the rest of the network forms a readout layer devoted to decode and interpret the high-dimensional dynamical state of the recurrent core. In that way, GRNs act as echo-state networks, providing the organism with a memory mechanism necessary to process temporal information.
Another important remark is that the training schemes used traditionally in the RC paradigm in the ANN context, such as ridge regression (Eq. 3), are not biologically realistic. To address this issue Gabalda-Sagarra, Carey, and Garcia-Ojalvo [11] studied whether an evolutionary process could successfully train the readouts. Their results showed that the output weights could indeed be tuned through an evolutionary algorithm, mimicking the way in which evolution would act upon the GRNs to enable computation.
IV Evolutionary aspects of recurrence-based computations
Besides the role of evolutionary processes in the feasible training of the output readouts in the RC framework, there are other relevant evolutionary aspects of RC that are worth highlighting here. In particular, adequate design properties are needed to produce operating reservoirs, as discussed by Seoane [51]. As for all biological strategies, it is expected that critical design decisions are influenced by evolutionary constraints. In the following we review the evolutionary constraints that underlie optimal operating reservoirs.
When RC was initially introduced, it was shown mathematically that the reservoir needs to have two fundamental attributes for the proper implementation of the paradigm [14]. Firstly, reservoir dynamics must be able to separate input features that are significantly different for a range of tasks. Secondly, these same dynamics should be able to generalize to new examples, thus effectively mapping similar inputs to proximate regions within the dynamical reservoir space.
Legenstein and Maass [52] analyzed how reservoir dynamics contributes to those two requirements, showing that separability and generalization are anticorrelated with each other: while recurrent networks with simple attractor dynamics have substantial generalization capabilities (because large classes of noisy inputs result in converging dynamics), they exhibit low separability. On the other hand, recurrent networks with chaotic dynamics have large separability but very low generalization capabilities (because chaotic systems are highly sensitive to initial conditions).
Consequently, to fulfill simultaneously the conditions of separability and generalization, recurrent networks need to operate in between ordered and chaotic dynamics, i.e. near the so-called edge of chaos. This refers to a critical state between ordered dynamics (where disturbances quickly die out into attractors) and chaotic dynamics (where disturbances are amplified). This property is particularly interesting given the evidence in the literature that cortical circuits are tuned to criticality [53, 54, 55].
In this sense, several authors have used hallmark indicators of criticality to design optimal reservoirs with enhanced performance. The measures employed include Lyapunov exponents [52, 56, 57, 58, 59] and Fisher information [60] of reservoir dynamics. Taken together, these studies highlight the advantages of criticality for information processing, and suggest powerful evolutionary constraints that should bring naturally occurring RC-based systems towards the edge of chaos.
V Final remarks
Living organisms are faced with the ongoing challenge of processing multiple simultaneous signals coming from their dynamical environment. In spite of the ubiquity of this requirement, a general conceptual framework of how biological systems process temporal information is still lacking. The intrinsically dynamical nature of recurrent networks allows in a natural way the integration of the present conditions faced by the living organism with its (recent) past, making the concept of recurrence-based computation an appealing initial step towards a conceptual model of temporal information processing in living beings.
In spite of their conceptual appeal, validating experimentally the role of recurrence-based computations in biological systems is still challenging. For example, while the existence of recurrent architectures in cortical brain regions has been examined in many works, as described above, studies linking systematically the dynamics of recurrent cortical networks with time-dependent tasks are still needed. To address this issue, cognitive systems should be evaluated simultaneously from the structural, dynamical, and functional viewpoints.
Model animals could be a suitable starting point for that goal. For instance, the fully mapped connectome of the roundworm C. elegans has been shown to be consistent with a network architecture that supports recurrence-based computations. When the dynamics of this network is compared with experimental observations, the system is seen to function in an input-driven chaotic regime that supports long-term memory via the phenomenon of generalized chaos synchronization [61]. Future interdisciplinary research should bridge the existing substantial literature on dynamical systems with our emerging knowledge of human and animal brains, together with our increasing understanding of the computational properties of recurrent neural networks. Given the fact that adaptability to changing environments is one of the main defining characteristics of living systems, the synergistic interaction between applied mathematics, neuroscience and machine learning discussed in this review can be expected to provide us with a general principle of biological information processing that has so far been elusive.
Acknowledgments
This work was supported by project PID2021-127311NB-I00, financed by the Spanish Ministry of Science and Innovation, the Spanish State Research Agency and FEDER (MICIN/AEI/10.13039/ 501100011033/FEDER). Financial support was also provided by the Maria de Maeztu Programme for Units of Excellence in R&D (Spanish State Research Agency, project CEX2018-000792-M), and by the ICREA Academia programme. M.S.V. is supported by a PhD fellowship from the Agència de Gestió d’Ajuts Universitaris i de Recerca (AGAUR) from the Generalitat de Catalunya (grant 2021-FI-B-00408).